Casey has 15 pounds of cashews that sell for $5.25 per pound. If peanuts sell for $2.50 per pound, how many pounds of peanuts should he add to cashews to obtain a mixturethat will sell for $3.75 per pound?
Answer by mananth(16946)

15 pounds of cashews $5.25 per pound.
peanuts $2.50 per pound, x pounds
mixture $3.75 per pound? x+15 pounds
...
15*5.25+2.50*x=3.75(x+15)
78.75+2.50x=3.75x+56.25
3.75x-2.50x=78.75-56.25
1.25x=12.50
x= 12.50/1.25
x= 10 pounds of peanuts
...
m.ananth@hotmail.ca